OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line (6,8)9x+y=3

OpenStudy (anonymous):

While parallel lines have the same slope, perpendicular have negative reciprocals: The first line can be written as y = -9x + 3 that means the perpendicular line must have a slope of 1/9 y = 1/9x + b substituting in the point: 8 = 1/9 (6) + b 8 = 2/3 + b b = 22/3 y = 1/9x + 22/3

OpenStudy (anonymous):

how about this one (4,-5)9x+8y=3??

OpenStudy (anonymous):

perpendicular or parallel?

OpenStudy (anonymous):

perpendicular

OpenStudy (anonymous):

First you need to write the first equation in y=mx+b form to find the slope: 9x+8y=3 8y=-9x+3 y=-9/8x+3/8 Then what did I say about the slope of a perpendicular line?

OpenStudy (anonymous):

it must have a slope

OpenStudy (anonymous):

but what must be true about that slope and the slope of the line it's perpendicular to?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

that's the slope intercept form, but the slope of a perpendicular line is always the negative reciprocal: The slope of the first line is -9/8 so the slope of the perpendicular line is 8/9 y=8/9x+b and then plug in the point: -5=8/9(4) + b -5 = 32/9 + b -77/9 = b y = 8/9x - 77/9